You plan to invest an amount of money in five-year certificate of deposit (CD) at your bank. The stated interest rate applied to the CD is 12 percent, compounded monthly. How much must you invest if you want the balance in the CD account to be $8,500 in five years? Please explain the formula.
Answer #1

We can use following formula to calculate the future value (FV) or worth of the investment

FV = PV * (1+i) ^n

Where FV = future value of deposits (CD) in five years =$8,500

PV or deposit or investment =?

And i= I/Y = 12% is the interest rate per annum but it is compounded monthly therefore monthly interest rate is 12%/12 = 1% per month

The time period n = 5 year or 5 years *12 = 60 months

Therefore,

$8,500 = PV * (1+1%) ^60

PMT = $4,678.82

You must have invest $4,678.82 if you want the balance in the CD account to be $8,500 in five years